SCI Computations

2020, latter half 2020, latter half



Gross purchases ₱280,000



Purchase returns and discounts 16% of the overall value



Beginning inventory Value of last semester’s ending inventory



Ending inventory 30% lower than the beginning inventory



Gross profit 60% based on cost of goods sold



Distribution expenses 10% of net sales



Administrative expenses 15% of gross profit



Income tax expense 30% of net income before taxes



Other comprehensive income 200% of income tax expense



Net income ? (6)



Cost of goods sold ? (7)



Other comprehensive income ? (8)



Comprehensive income ? (9)



Income tax expenses ? (10)
1
Expert's answer
2021-09-23T09:14:10-0400

Solution:

Calculate the cost of goods sold:

Cost of goods sold = Beginning inventory + Net purchases – Ending inventory

Beginning inventory = 106,600

Gross Purchases = 280,000

Net purchases = Gross purchases – Purchase returns & discounts

Purchase returns & discounts = 16"\%" of gross purchases value = 280,000 "\times 16\%" = 44,800

Net purchases = 280,000 – 44,800 = 235,200

Ending inventory = 30"\%" lower than the beginning inventory = 106,600 "\times 70\%" = 74,620

Cost of goods sold = 106,600 + 235,200 – 74,620 = 267,180

 

Gross profit = 60"\%" of Cost of goods sold = 267,180 "\times 60\%" = 160,308

Gross profit = Net sales – Cost of goods sold

Net sales = Cost of goods sold + Gross profit

Net sales = 267,180 + 160,308 = 427,488

Distribution expenses = 10"\%" of net sales = 427,488 "\times 10\%" = 42,748.80

Administrative expenses = 15"\%" of gross profit = 160,308"\times 15\%" = 24,046.20

Net income before taxes = Gross profit – Distribution expenses – Administrative expenses

Net income before taxes = 160,308 – 42,748.80 – 24,046.20 = 93,513

Income tax expense = 30"\%" of Net income before taxes = 93,513"\times 30\%" = 28,053.90

Net income = Net income before taxes – Income tax expense = 93,513 – 28,053.90 = 64,459.10

Other comprehensive income = 200"\%" of income tax expense = 28,053.90 "\times 200\%" = 56,107.80

Comprehensive income = Net income + Other comprehensive income

Comprehensive income = 64,459.10 + 56,107.80 = 121,566.90


Net income (6) = ₱64,459.10

 

Cost of goods sold (7) = ₱267,180

 

Other comprehensive income (8) = ₱56,107.80

 

Comprehensive income (9) = ₱121,566.90

 

Income tax expenses (10) = ₱28,053.90